Describe a particle’s motion when t is greater or equal to 0. Given x=4 - 27t + t^3
v=3t^2 - 27
a= 6t
Answer #1

slope of dv/dt is 6t which is positive for t>0 so there is positive acceleration and velocity is increasing for t >0 . at t=0 ,particle is moving negative direction from initial position x=4 with v(0) =3*0-27 = -27m/s and initial position x= 4-0-0 = 4m .